Tom read a book from page 1 to page 59 without skipping any page.What is the sum of all the page number?
1
Expert's answer
2011-10-13T11:35:14-0400
we must find sum 1+2+3+4+....+58+59
it is a sum of arithmetic progression
with difference d=1 so S=n(a1+an)/2, where n is total number, a1, an - fisrt and
n-th summand so we have S=60*(1+59)/2=60*30=1800